40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1625     Entry time: Thu Apr 7 02:00:08 2016
Author: Antonio 
Type: DailyProgress 
Category: PMC 
Subject: PMC locking 

Following the previous elog entry today we succeeded in locking the PMC.

 

As a reminder this is only a temporary setup. The optical path need to be re-design in order to allow

the implementation of the PMC together with mode matching lenses.  

This is the list of things we did today:

  • Setup RFPD on the table for the PMC
  • Added the EOM on the path (may be temporary, as the mount is not ready)
  • Connected the LO to EOM and the PMC card
  • Checked Error signal
  • Locked the cavity
ELOG V3.1.3-